OpenStudy (anonymous):

Use the arc length formula and the given information to find r. s = 16 cm, θ = 48°; r = ?

OpenStudy (anonymous):

\[s = \frac{\theta}{r}\] So: \[r = \frac{\theta}{s}\]

OpenStudy (anonymous):

Find r from here.. \(\theta = 48\) and \(s = 16\) here..

OpenStudy (anonymous):

48/16=3 so r =3 ? :)

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

thank you very much :)

OpenStudy (anonymous):

Welcome dear..